There are twenty multiple-choice questions on an exam, each having responses a, b, c, or d. Each question is worth 5 points and only one response per question is correct. Suppose a student guesses the answer to each question, and her guesses from question to question are independent. On average, this strategy will result in a mean score of

Select one:

a. 5.

b. 25.

c. 50.
Answer #1

here this is binomial distribution with parameter n=20 and p=probability of guessing it right ==1/4

hecnce expected correct answer =np=20*1/4=5

therefore expected mean score =5*5=25

option B is correct